OpenStudy (anonymous):

Twice a number d? Isn't that d^2?

OpenStudy (angela210793):

twice a no. d = 2d

OpenStudy (anonymous):

Twice a number is 2d A number squared = d^2

OpenStudy (anonymous):

twice=2 times if 2d=d^2 d=5 2(5)=(5)^2 10=25 which is not true so twice a number is 2d not d^2